OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (amoodarya):

hint : write sin(2x)=2sinx cos x it will be easy to solve

OpenStudy (anonymous):

Thanks (: But, how do you find solutions? I'm not sure where to begin.

OpenStudy (anonymous):

@amoodarya I wrote out the equation, but am stuck where to begin.

OpenStudy (amoodarya):

let's do it step by step

OpenStudy (anonymous):

Ok (:

OpenStudy (amoodarya):

a question ? is it sin(2x) or sin^2(x)

OpenStudy (anonymous):

The equation is: 7sin^2(x)-14sinx+2=-5

OpenStudy (amoodarya):

\[7\sin^2(x)-14sinx+2=-5\\7\sin^2 x-14 sinx +2+5=0\\7\sin^2 x-14 sinx +7=0\\ \div 7\\sin^2x-2sinx+1=0\\(sinx-1)^2=0\\\]

OpenStudy (anonymous):

So, to find the solutions, I just solve for x in the equation: (sinx-1)^2=0?

OpenStudy (amoodarya):

yes . can you go on ?

OpenStudy (anonymous):

Yes, I've got it now, thanks!! (:

OpenStudy (anonymous):

Wait...actually. Could you explain how to find the solutions? I'm still unsure what to do about finding the solutions.

OpenStudy (anonymous):

@amoodarya ^^ (:

OpenStudy (amoodarya):



OpenStudy (anonymous):

So, would my answers be pi/2 and 3pi/2?

OpenStudy (anonymous):

^ @amoodarya